OpenStudy (anonymous):

PLEASE HELP ME!!!!!!!! WILL FAN+MEDAL!! Which table shows a proportional relationship?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (butterfield1215):

i would say a

OpenStudy (anonymous):

let me give you a little hint....proportions are nothing but equivalent fractions

OpenStudy (anonymous):

thanks! anyone else think it is a?

OpenStudy (butterfield1215):

what do you think it is

OpenStudy (anonymous):

its not A look for equivalent fractions

OpenStudy (anonymous):

I'm not really sure?

OpenStudy (anonymous):

what about D?

OpenStudy (butterfield1215):

sorry c i mixed that one up

OpenStudy (anonymous):

is it D @kelliegirl33 ?

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

look at C 2/-4 = -1/2 3/-6 = -1/2 5/-10 = -1/2 6/-12 = -1/2 it is C.....nothing but equivalent fractions

OpenStudy (anonymous):

well, thanks so much!!

OpenStudy (anonymous):

sure thing :)